MICQRC(5) MICQRC(5) ˜ /.micq/micqrc−mICQ NOTE-THISFILEHASNOTYETBEENUPDATEDFORmICQ0.4.10. mICQ.,[],-. CONNECTIONS ICQ.,,:ICQ,--. typetype[auto] type,serverICQ,peer,--.auto,.,,,,,UIN’,,UIN’,. version .,5,8,--6. server ,.icq.icq.com.1--1. port,,1--1. uinUIN UIN,UIN.1--1. password[] ,.1--1.,. status ,.: 0Online 1 5 17 19 32 256 --,: 1.. 2,SOCKS. 4.. GENERAL SOCKS5,,-. s5use 01. s5host socks5. s5port socks5. s5auth 01. RU mICQ 1 MICQRC(5) MICQRC(5) s5name socks5. s5pass socks5. : verbose . autoaway .0. Screenwidth .,0,80.,... setdelbson|off on,deletebackspace.off,. setrussianon|off koi8-r/ucp1251.,on. setjapaneseon|off Shift-JISEUC. setfunnyon|off "funny". setcoloron|off . sethermiton|off on,,,-. setlogon|off . setlogonoffon|off ,online/offline. setautoon|off . setuinprompton|off ,. setlinebreak ,simple,break,indentsmart,. settabssimple|cycle|cycleall,.mICQsimple,msg,.cycleonline-.msg. cycleallcycle,,online. chatnr -nr.-1,49mICQ(). ,-: colornr nr. colorusecolor coloruse.usenone,server,client,message,contact,sent,ack,error,debug incoming,colorblack,red,green,yellow,blue,magenta,cyan,white,none,bold (bold,,),,. RU mICQ 2 MICQRC(5) MICQRC(5) logplace| ,,,.,/. soundon|off|cmd .onoff,. soundonlineon|off|cmd ,,online. soundofflineon|off|cmd ,,offline. receivescriptcmd ,.. auto.away,na,dnd,occ,inv. STRINGS . alteroldnew oldnew.,new.,micq(7).. CONTACTS -. [*][˜ ][ˆ ]uin UIN’uin.*,.˜ ,offline.ˆ ,.uin,,,. :-. micq(1),micq(7) JamesMorrison<ja2morrison@student.math.uwaterloo.ca>.Ru¨ digerKuhlmann <micq@ruediger-kuhlmann.de>."Pest"<pezt@mail.ru>. RU mICQ 3